# Data Augmentation for Supervised Graph Outlier Detection with Latent Diffusion Models

### v2

@@ IV-A2 Baselines
IV-A3 Metrics
We follow the extensive literature in graph outlier detection [5, 6] to comprehensively evaluate the performance of graph outlier detectors with three metrics that are robust to class imbalance: Receiver Operating Characteristic-Area Under Curve (AUC), Average Precision (AP), and Recall@k (Rec), where the value of k is set to the number of actual outliers present in the dataset.
IV-A4 Implementation Details
We modified GADBench to benchmark outlier detection performance.
Environment. The key libraries and their versions used in our experiments are as follows: Python 3.9, CUDA 11.8, PyTorch 2.0.1 [42], PyTorch Geometric 2.4.0 [43], DGL 1.1.2 [44], and PyGOD 1.0.0 [4].
TABLE III: Performance comparison in AUC, AP, and Rec (%) on four datasets.
Dataset Weibo Tolokers Questions Elliptic
Metric AUC AP Rec AUC AP Rec AUC AP Rec AUC AP Rec
GCN 98.11 93.48 89.34 74.69 42.88 42.06 69.81 12.54 16.99 82.68 22.23 27.61
SGC 98.66 92.46 87.90 70.67 38.03 35.98 69.88 10.13 15.62 73.02 11.44 9.14
GIN 97.47 92.67 87.90 74.05 36.57 36.76 67.76 12.30 18.36 84.38 29.66 35.64
GraphSAGE 96.54 89.25 86.17 79.42 48.65 46.42 71.69 17.63 21.10 85.31 37.52 36.20
GAT 94.08 90.25 86.74 77.26 43.14 43.30 70.33 14.51 17.26 84.42 23.43 27.42
GT 97.06 91.44 87.03 79.24 46.22 46.57 70.83 16.14 20.27 87.14 29.91 38.97
[…]
DCI 93.90 87.78 83.86 75.98 39.85 40.19 67.95 14.58 19.18 81.93 27.63 33.15
[…]
GATSep 96.72 91.55 89.05 79.63 46.08 46.73 69.96 15.98 19.18 83.89 21.46 21.15
BernNet 93.85 88.00 85.30 76.20 42.20 42.21 70.80 16.04 17.53 82.01 20.52 23.55
AMNet 95.88 89.74 85.59 75.83 42.66 41.90 69.71 17.02 19.18 80.06 16.73 17.17
BWGNN 98.29 92.72 84.73 80.15 49.65 47.35 69.47 16.24 18.63 84.32 22.56 26.50
GHRN 97.21 92.67 88.18 79.80 49.50 48.29 68.24 16.24 18.63 85.36 24.01 30.29
DAGAD 98.54 83.36 90.78 77.69 33.94 44.39 71.21 6.88 20.55 85.62 26.18 40.54
[…]
Hyperparameters. GODM is implemented with default hyperparameters. For Variational Encoder, we use one layer GraphSAGE for $\text{GNN}_{\text{shared}}$, $\text{GNN}_{\bm{\mu}}$, and $\text{GNN}_{\bm{\sigma}}$, respectively. We set the hidden dimension to the largest power of 2, which is no greater than the feature dimension of the dataset divided by 2. The weights of reconstruction loss $\omega_{\mathbf{X}}$, $\omega_{\mathcal{E}}$, $\omega_{\mathbf{t}}$, and $\omega_{\mathbf{p}}$ are 1, 0.5, 1, and 0.3, respectively. The weight for KL-divergence $\beta$ is 0.001. For the diffusion model, we use the dimension twice as the hidden dimension in VAE. We use five layers of MLP as the denoising function following the detailed design in [22]. For training, we adopt the Adam optimizer with a learning rate of 0.001 without weight decay to train the VAE and the diffusion model for 100 epochs, respectively, and apply early stopping with a patience of 50. The negative sampling ratio is 1, and the approximate graph partition size is 2048. For inference, we use 50 steps of diffusion and generate the same amount of synthetic outliers as the number of real organic outliers in the training set. For all the hyperparameters in graph outlier detection, we apply the default setting in GADBench. We report the graph outlier detection performance of GODM with the optimal downstream graph outlier detector.
Open-Source Package. To enhance accessibility, we encapsulate our data augmentation into user-friendly API and make our code a plug-and-play package. This package is built upon PyTorch and PyTorch Geometric (PyG) frameworks. It accepts a PyG data object as input and returns the augmented graph as output. By inheriting the transforms.BaseTransform in PyG, we enable users to apply GODM just like any other transformations supported in PyG, e.g., ToUndirected. We are also able to customize hyperparameters during initialization. Code Demo 1 gives a brief example of API usage.
1from pygod.utils import load_data
2data = load_data(’weibo’) # load data
@@ IV-A4 Implementation Details
4from godm import GODM # import GODM
5godm = GODM(lr=0.004) # init. GODM
6aug_data = godm(data) # augment data
@@
8detector(aug_data) # train on data
@@
For distribution, we release our package on the Python Package Index (PyPI). Users can easily install GODM by executing pip install godm in a command line.
IV-B Performamce on Graph Outlier Detection
TABLE IV: GODM’s performance improvements on graph outlier detection algorithms in AUC, AP, and Recall (%) on Tolokers.
AUC AP Rec
GCN 75.45 (+0.76) 44.17 (+1.29) 44.24 (+2.18)
SGC 72.73 (+2.06) 39.75 (+1.72) 38.01 (+2.02)
GIN 74.83 (+0.78) 38.54 (+1.96) 38.32 (+1.56)
GraphSAGE 81.65 (+2.23) 52.53 (+3.87) 50.93 (+4.52)
GAT 82.18 (+4.91) 51.13 (+7.99) 50.78 (+7.48)
GT 82.73 (+3.49) 51.73 (+5.51) 50.93 (+4.36)
GAS 76.96 (+0.05) 45.48 (-1.87) 43.15 (-1.87)
DCI 73.75 (-2.23) 37.52 (-2.33) 37.54 (-2.65)
PCGNN 73.65 (+1.47) 38.42 (+0.90) 38.01 (+1.25)
GATSep 83.46 (+3.83) 52.95 (+6.87) 52.80 (+6.07)
BernNet 76.53 (+0.33) 44.04 (+1.83) 42.21 (+0.00)
AMNet 76.67 (+0.84) 44.31 (+1.65) 43.30 (+1.40)
BWGNN 82.10 (+1.95) 51.94 (+2.29) 51.25 (+3.89)
GHRN 82.05 (+2.25) 51.84 (+2.34) 51.56 (+3.27)
We start by examining how effective GODM is in the task of graph outlier detection on homogeneous graphs. We report graph outlier detection performance on four homogeneous datasets in terms of AUC, AP, and Rec of different algorithms in Table III. In the table, the highest score in each metric is marked in bold, while the second highest score is underlined. For DAGAD, we report the optimal performance of its two implementations with GCN and GAT. Although DAGAD shows superior performance compared to the standalone GCN and GAT, it fails to surpass more sophisticated detectors. On the other hand, GODM and GOVAE together demonstrate significant enhancements in graph outlier detection performance across all datasets and metrics. Notably, on the Weibo dataset, where the performance metric is already relatively high (AUC of 98.66 for SGC), GODM further elevates the performance, achieving an AUC of 99.57. A particularly noteworthy improvement is observed on Elliptic, where GODM enhances the Rec by over 20% (an increase of 9.14) compared to GAS, which is the second-best performing algorithm. Although GOVAE exhibits competitive performance, its limited generation ability results in generally weaker performance compared to GODM, especially on Elliptic.
Further investigations were carried out to understand the extent of improvement GODM could bring to various graph outlier detection algorithms, specifically on Tolokers. We integrate GODM on the top of all graph outlier detection algorithms and report the performance changes in terms of AUC, AP, and Rec in Table IV. The absolute performance is reported out of brackets, and the relative change is reported in the brackets. We can observe that GODM enhances graph outlier detection performance for most of the algorithms. GODM increases the AP of GAT by 7.99. For DCI and GAS, the only two algorithms that minorly decrease the performance, this could be attributed to suboptimal hyperparameter settings for these specific algorithms or the dataset. We believe the performance can be improved by careful hyperparameter tuning.
TABLE V: Performance comparison in AUC, AP, and Rec (%) on DGraph.
AUC AP Rec
HGT 72.89 3.13 5.20
w/o Feature 75.10 3.64 5.59
w/o Edge 74.36 3.52 5.76
w/o Time 74.23 3.29 4.73
w/o Type 75.09 3.62 5.98
GOVAE 75.36 3.58 5.54
GODM 75.80 3.66 6.19
Moreover, in Table V, we benchmark the graph outlier detector performance on DGraph, a large-scale dataset with different types of edges and temporal information. We include HGT along with variants of GODM. For the w/o Feature, we randomly sample the node feature from the Gaussian distribution to replace the generated node feature in the synthetic graph. For the w/o Edge, all generated edges are substituted with random edges. For w/o Time, we give uniformly sampled timestamps to replace the generated timestamps. For the w/o Type, we use a random type sampled from a uniform distribution for all generated edges. According to Table V, GODM demonstrates enhanced performance across all metrics when compared to HGT. Furthermore, masking any generated information leads to a notable decline in performance, which underscores the significance of each aspect of the generated data. While GOVAE achieves commendable results, its performance still falls short of GODM’s, highlighting the superior efficacy of Latent Diffusion Model.
IV-C Generation Quality
To assess the quality of generated data, we conduct a case study on DGraph to compare real data and synthetic data from both GOVAE and GODM. The results are shown in Fig. 3. Specifically, Fig. 3 (a) and (b) illustrate the distribution density of a single dimension in node features. From the figures, we can see that the distributions of GODM generated node features are close to the complex distributions of real data, while GOVAE can only generate Gaussian distributions, showcasing the generation ability of Latent Diffusion Model. In addition, to evaluate generated edges, Fig. 3 (c) shows the frequency across eleven different edge types. The edge type generated by GODM has a more similar distribution to real data compared to GOVAE, particularly on the 10th edge type. It further underscores the superiority of GODM over GOVAE.
Figure 3: Visualization of single node feature density and edge type frequency of real data and synthetic data from GOVAE and GODM on DGraph.
IV-D Efficiency Study
We further evaluate the efficiency of GODM and its variants in terms of time and memory on graphs of various sizes, including Weibo, Elliptic, and DGraph. In order to demonstrate the efficiency of GODM, we remove graph clustering (w/o GC), negative sampling (w/o NS), and EDM (w/o EDM) for evaluation. To simulate the real-world application, we measure the running time of unit epoch training for VAE and Latent Diffusion Model plus the inference process. For memory, we present the maximum active GPU memory usage within the whole data augmentation, as the GPU memory constraint is usually the bottleneck for machine learning systems. The results are shown in Fig. 4.
Figure 4: Running time and GPU memory of GODM and its variants.
From Fig. 4, we can observe that employing a graph clustering algorithm significantly diminishes GPU memory utilization. Specifically, on DGraph, w/o GC results in a memory usage that is over 30$\times$ higher than GODM. On the other hand, graph clustering turns full-batch training into mini-batch training, leading to an anticipated increase in running time. Additionally, negative sampling reduces both the running time and the memory usage, as the number of edges in the training set reduces from $O(n^{2})$ to $O(|\mathcal{E}|)$. While EDM maintains consistent memory consumption, it contributes to time savings by minimizing the number of sampling steps.
V Related Works
This section briefly summarizes the previous related works in three key areas, including graph outlier detection, class imbalance, and graph generative models.
V-A Graph Outlier Detection
Graph outlier detection is a vital task in data mining and machine learning, aiming to identify anomalous structures (e.g., nodes, edges, subgraphs) in graphs. Due to its structural simplicity and broader applications, most research focuses on node-level graph outlier detection and can easily convert to edge level and graph level task [4]. Many studies focus on unsupervised settings, detecting outliers in the graph only based on the graph data itself without any ground truth labels [45, 46, 47]. However, such unsupervised approaches may not align well with scenarios necessitating the detection of specific outlier types containing domain-specific knowledge. In this case, (semi-)supervised graph outlier detectors, which can learn ground truth labels, are better options. [33] apply attention mechanism to detect spam review. [11, 48] alleviate the camouflage issue in fraud detection by message passing in selected neighborhoods. [35] further improves the neighbor selection in fraud detection. [49] builds the outlier detection graph via the KNN algorithm. [34] decouples the representation learning and anomaly detection, and [50] focuses on heterophily in graph outlier detection. Some other works detect outliers from a spectral perspective. [36] learns arbitrary graph spectral filters via Bernstein approximation, and [38] is capable of discerning both low-frequency and high-frequency signals, thereby adaptively integrating signals across a spectrum of frequencies. In [39], the Beta kernel is employed to detect anomalies at higher frequencies using flexible and localized band-pass filters. [40] addresses heterophily in graph outlier detection from a graph spectrum perspective.
V-B Class Imbalance
Despite the fruitful literature on graph outlier detection, most of methods underrate a pivotal challenge: the class imbalance in ground truth labels. These methods typically mitigate this imbalance by primitive techniques. GADBench uses reweighting [6], assigning a higher weight to outliers in the loss function. Some works adopt upsampling to replicate outliers, which is mathematically equivalent to reweighting. [11] applies downsampling, reducing the number of normal instances and losing a large amount of valuable supervision information from normal instances. However, these tricks may not sufficiently address the underlying complexities associated with imbalanced datasets. Consequently, there is a compelling need for more sophisticated techniques that can effectively address this class imbalance to ensure the quality of graph outlier detection. [35] balances the neighborhood distribution by reducing edges between different classes and adding edges between the same classes but does not change the overall number of positive and negative nodes. [41] adopts random perturbation for data augmentation to generate more samples. Other general graph machine learning studies apply interpolation-based data augmentation methods in latent space [51]. However, these methods are not specifically designed for graph outlier detection, and interpolation is too naive to generate helpful instances. This paper seeks to build upon these foundational works, proposing a novel graph data augmentation algorithm with latent diffusion models that not only mitigate class imbalance effectively but are also scalable and efficient on large graphs, ensuring broader applicability in real-world graph outlier detection.
V-C Graph Generative Models
Recent advances in graph generative models have catalyzed a significant body of research, focusing primarily on the synthesis of realistic and structurally coherent graphs. [52] proposes VGAE, leveraging variational autoencoder frameworks to learn latent representations of graph structures for generation purposes. [53] introduces GraphRNN, generating nodes and edges autoregressively with two recurrent neural networks. [54] applies a generative adversarial networks framework to generate graphs via random walks. [55] introduces the normalizing flow method for molecular graph generation. With the recent proliferation of works in image generation with diffusion models, some efforts have extended the generation power of diffusion models to graph-structured data. [13] proposes a score-based generative modeling of graphs via the system of stochastic differential equations, while [14] uses a discrete diffusion model to generate graphs. However, the potential of graph generative models to address the class imbalance in graph outlier detection remains largely underexplored. We propose to generate some synthetic outliers to mitigate this class imbalance in graph outlier detection.
VI Conclusion
In this paper, we introduce GODM, a novel method for data augmentation in supervised graph outlier detection. GODM is designed to address the significant challenge of class imbalance, a pervasive issue in graph outlier detection, where the scarcity of outlier instances compared to normal instances often leads to suboptimal model performance. Our method leverages the power of generative models, particularly latent diffusion models, to synthesize fidelity graph data that is statistically similar to real outliers. GODM consists of three key components: Variational Encoder, Latent Diffusion Model, and Graph Generator. Variational Encoder maps heterogeneous information in the input graph into a unified latent space. Latent Diffusion Model learns the latent space distribution of real organic data through iterative denoising. Graph Generator then synthesizes graph data from this latent space embedding, ensuring statistical similarity to real data. Our extensive experiments across multiple datasets demonstrate the effectiveness and efficiency of GODM. Note that GODM is model agnostic, which means it can be flexibly integrated with different downstream graph outlier detectors. We encapsulate GODM into a plug-and-play package, making it accessible for broader use in the researcher and practitioner communities.
@@ VI Conclusion
In future research, several directions can be explored to enhance GODM further and expand its applicability in the domain of graph outlier detection:
@@
  • Collaboration with outlier detection. Currently, GODM is model agnostic, meaning the training of GODM and downstream outlier detection tasks are separate. Integrating the supervision signals from both the generation task and the downstream outlier detection in a unified training framework could be mutually beneficial. Self-training offers a promising paradigm for this integration [57].
  • Enhancing generative models. Although the current diffusion model in GODM has shown impressive capabilities, there is potential for further improvement. This could be achieved by employing more expressive denoising functions, such as transformers [58], or by exploring alternative generative models like energy-based models [59] and or normalizing flows [60], which might offer different advantages over diffusion models.
Acknowledgment
This work is supported by NSF under grant III-2106758. We sincerely appreciate reviewers in advance for their significant effort and invaluable feedback during the review process.

### v3

@@ G.3 Data Augmentation
Appendix H Description of Metrics
AUC (Receiver Operating Characteristic-Area Under Curve): AUC quantifies the area beneath the Receiver Operating Characteristic Curve, which is constructed by plotting the true positive rate against the false positive rate across varied determined threshold levels. An AUC of 1 indicates flawless predictive accuracy, whereas an AUC of 0.5 indicates an absence of discriminative power, equivalent to random guessing. This metric is favored over accuracy for evaluating outlier detection models due to its robustness against class imbalance prevalent within the class distributions.
AP (Average Precision): AP offers a comprehensive summary of the precision-recall curve, represented as the weighted average of precision values attained at each threshold, utilizing the increment in recall from the preceding threshold as the weight. This metric offers a balance between recall and precision, with a higher AP signifying a lower rate of both false-positive rate (FPR) and false-negative rate (FNR). For outlier detection applications, e.g., fraud detection, both FPR and FNR are critical, as misclassification could either cause financial loss or harm normal user experience.
Rec (Recall@k): Considering the minority of outliers relative to the abundance of normal instances in datasets, Recall@k is proposed as a measure of how well the detectors rank outliers over the normal samples. The value of k is set to the number of actual outliers present in the dataset. Recall@k is then determined by the ratio of the true outliers within the top k-ranked samples to k. A maximal Recall@k value of 1 means a model perfectly ranks all outliers over normal samples, while minimal value of 0 means none of the outliers is ranked at top k.
Appendix I API Demo
1 from pygod.utils import load_data # import data
2 data = load_data(’weibo’) # load weibo data
@@ Appendix I API Demo
4 from godm import GODM # import GODM
5 godm = GODM(lr=0.004) # initialize GODM
6 aug_data = godm(data) # augment data
@@
8 detector(aug_data) # train on data
@@
Appendix J Future Directions
@@
For future research, here are several promising directions we may explore to further enhance GODM and expand its applicability in the domain of graph outlier detection:
@@
  • Collaboration with outlier detection. Currently, GODM is model agnostic, meaning the training of GODM and downstream outlier detection tasks are asynchronous. On the other hand, integrating the supervision signals from both the generation task and the downstream outlier detection in a unified training framework could be mutually beneficial. Self-training offers a promising paradigm for this integration of the supervision signals [59].
  • Enhancing generative models. Although the current diffusion model in GODM has shown impressive capabilities, there are opportunities for further improvement. This could be achieved by employing more expressive denoising functions, such as transformers [60], or by exploring alternative generative models like energy-based models [61] and or normalizing flows [62], which might offer different advantages over diffusion models.